the tens digit of a 2 digit number is 3 times the units digit. If the digits are reveresed, the new number is 36 less then the original number. find the original number?
Answer by checkley75(3666)

X=3Y
10X+Y=10Y+X+36
10(3Y)+Y=10Y+3Y+36
30Y+Y=13Y+36
31Y-13Y=36
18Y=36
Y=36
18
Y=2 ANSWER FOR THE ORIGINAL UNITS DIGIT.
THUS THE ORIGINAL TENS DIGIT IS 3*2=6
PROOF:
62-26=36